OpenStudy (anonymous):

6a^2+5a-4 factor..someone plz?

jimthompson5910 (jim_thompson5910):

6a^2+5a-4 6a^2+8a-3a-4 (6a^2+8a)+(-3a-4) Keep going to completely factor

OpenStudy (anonymous):

is (12a+20a) (-3a-7) the final answer??

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

Then wat is it?? :S

jimthompson5910 (jim_thompson5910):

what can you factor out of 6a^2+8a

OpenStudy (anonymous):

a 2??

jimthompson5910 (jim_thompson5910):

close, more like 2a

jimthompson5910 (jim_thompson5910):

6a^2+8a = 2a(3a + 4)

OpenStudy (anonymous):

(3a-4)(a+1) a=4/3 or a=-1

jimthompson5910 (jim_thompson5910):

now, is it possible to write (-3a-4) in the form k(3a+4) ?

OpenStudy (anonymous):

i dont know.... you tell me?There are no k's so y wud u add one??? :s

jimthompson5910 (jim_thompson5910):

k can be any number or expression, think of it as a variable

OpenStudy (anonymous):

ohh ok

jimthompson5910 (jim_thompson5910):

ok let me ask this is it possible to factor (-3a-4) to get (3a+4) times SOMETHING where that SOMETHING is an expression or number ??

OpenStudy (anonymous):

huh???

jimthompson5910 (jim_thompson5910):

is it possible to turn (-3a-4) into (3a+4) ?

OpenStudy (campbell_st):

here is a really simple method for any difficult factorisation if the quadratic is Your question ax^2 + bx + c = 0 6a^2 + 5a - 4 = 0 1. multiply a and c 6 x -4 = -24 2. Find the factors of a*c that add to b factors -3 + 8 = 5 next step up a fraction with a as the denominator the numerator is in the form (ax + factor 1)(ax + factor 2) for your question  this meothod works everytime... you don't need to split the middle term...